OpenStudy (anonymous):

How would I find the answer to 12*3.14?

OpenStudy (anonymous):

12 x 3.14 I can never remember how to multiply decimals

OpenStudy (jiteshmeghwal9):

\[\Huge{12 \times \frac{314}{100}}\]

OpenStudy (accessdenied):

Basically, you just multiply 12 x 314 without the decimals as normal, and then you move the decimal however many places total from the original two numbers. Like, since you had two decimal places to the left in 314, you move the end result of 12x314's decimal place to the left two places as well. It makes sense because decimals are usually caused by fractions of powers of 10. 3.14 is the same as 314/100, and if you multiply 12x3.14 = 12x314/100, you multiply the numerators.

OpenStudy (jiteshmeghwal9):

can u solve this ?

OpenStudy (jiteshmeghwal9):

now there is nt any decimal problem :)

OpenStudy (jiteshmeghwal9):

aliter method :- simply multiply 314 to 12 and then put decimal point before 2 digits .

OpenStudy (anonymous):

12*3.14= 37.68 Is this correct?

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (jiteshmeghwal9):

right @mayankdevnani ?

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (accessdenied):

You're welcome! :)